OpenStudy (anonymous):

What's the difference between f(x)^2 and f(x^2)

OpenStudy (jamesj):

Consider the function f(x) = x + 1. Then \[ f(x^2) = x^2 + 1 \] but \[ f(x)^2 = (x+1)^2 = x^2 + 2x + 1 \] Hence they are not equal in general.

OpenStudy (aravindg):

hey